OpenStudy (anonymous):

What is the solution to the inequality: 2x+4 >and equal to 7

OpenStudy (anonymous):

3 <= x <= 6

OpenStudy (anonymous):

the choices are a. x>and equal to 3/2 b. x< equal to 3/2 c. x> equal to 11/2 d. x<equal to 11/2

OpenStudy (anonymous):

so it wouldd be b

OpenStudy (anonymous):

ok how u figure it out

OpenStudy (anonymous):

how do i figure this out

OpenStudy (anonymous):

\[ 2x+4 \ge 7\\ 2x \ge 7 -4=3\\ x \ge \frac 3 2 \] a. is the answer